# Optimal Dynamic Fees in Automated Market Makers

### v1

@@ Abstract
Automated Market Makers (AMMs) are emerging as a popular decentralised trading platform. In this work, we determine the optimal dynamic fees in a constant function market maker. We find approximate closed-form solutions to the control problem and study the optimal fee structure. We find that there are two distinct fee regimes: one in which the AMM imposes higher fees to deter arbitrageurs, and another where fees are lowered to increase volatility and attract noise traders. Our results also show that dynamic fees that are linear in inventory and are sensitive to changes in the external price are a good approximation of the optimal fee structure and thus constitute suitable candidates when designing fees for AMMs.
@@ 1 Introduction
The early works of Capponi and Jia (2021), Bartoletti et al. (2021), Angeris and Chitra (2020), Angeris et al. (2021), and Angeris et al. (2023) first explain how these exchanges operate and introduce the problem of liquidity provision in AMMs. As shown by Cartea et al. (2023, 2024b) and Milionis et al. (2022), liquidity provision in the absence of fees leads to losses for liquidity providers. This issue, known as impermanent loss (IL), occurs because liquidity provision in an AMM is passive; when the prices of risky assets move, the liquidity providers trade at worse than market prices.
To address the impermanent loss problem, a significant part of the research has focused on improving the design of AMMs to make liquidity provision less exposed to arbitrageurs. Some works in this direction are Adams et al. (2025), where they set up an auction-based AMM in which LPs are rewarded by a so-called manager of the pool who can profit from arbitrage opportunities; Goyal et al. (2023), where they translate LPs’ beliefs about future asset valuations into an optimal choice of a trading function; Cartea et al. (2024a), where they introduce a more general version of constant function market makers (CFMs) called decentralized liquidity pools (DLPs), in which liquidity providers prevent predictable losses by choosing particular quote functions, impact functions, and monitoring external information; and Bergault et al. (2024b), where they design an oracle-based AMM in which price discovery does not rely solely on liquidity takers, and the pricing function incorporates external information about the current market price of the risky asset. Recently Capponi and Zhu (2025) addressed the problem of IL by determining an optimal exit strategy for the liquidity provider. Another interesting contribution is the work of Fukasawa et al. (2023), where the authors hedge impermanent loss in a constant function market maker using a weighted variance swap.
@@
As shown by Milionis et al. (2024), impermanent loss can be partially or fully hedged through the collection of fees. The problem of determining optimal fees for AMMs has been an important issue since the early stages of research in this area. Evans et al. (2021) focused on how to find optimal fees for geometric mean market makers (G3Ms) while Fritsch (2021) studied how to determine optimal fees in a setting of multiple liquidity pools competing with each other for trade volume. However, both works only address the problem of finding constant static fees. Sabate-Vidales and Šiška (2023) address the impermanent loss problem by setting dynamic fees although no optimality is established. Hasbrouck et al. (2022) found how higher fees on decentralised exchanges (DEX) can increase trading volume because higher fees attract more liquidity providers and boost the inventory in the DEX. Recently, Aqsha et al. (2025) solve a stochastic leader-follower game between the AMM and the liquidity providers in the venue. They find that under the equilibrium contract, the liquidity providers have incentives to add liquidity to the pool only when higher liquidity on average attracts more noise trading. In their work, the fee that the venue charges is proportional to the trade size.
@@
In this study, we address the problem of determining optimal dynamic fees for AMMs. As argued in Cao et al. (2023), the fees should adjust dynamically in order to reflect market conditions. With the recent introduction of Uniswap v4, the implementation of dynamic fees is now possible. Setting optimal fees is crucial because fees are the primary incentive for liquidity providers to engage with the AMM as they help offset losses from impermanent loss. As a result, a platform that optimizes fee collection can better compensate liquidity providers and, in turn, strengthen its overall market position. Inspired by the work on market making by Avellaneda and Stoikov (2008), [^1] For a comprehensive exposition of the market making problem and its generalisation see the monographs by Cartea et al. (2015) and Gueant (2016). we formulate a stochastic control problem in which the venue aims to maximize the total fees collected in a liquidity pool consisting of two assets: a risky asset $Y$ and a riskless asset $X$. The control processes are $\mathfrak{p}$ and $\mathfrak{m}$, representing the fees charged for selling and buying asset $Y$, respectively. [^2] We choose $\mathfrak{p}$ for plus and $\mathfrak{m}$ for minus because the quantity of asset $Y$ in the pool increases after a trade that pays $\mathfrak{p}$ and decreases after a trade that pays $\mathfrak{m}$.
These fees influence the exchange rate faced by liquidity takers: buying $Y$ requires paying more of asset $X$, while selling $Y$ yields less of asset $X$. This, in turn, affects the order flow. We model sell and buy orders using two point processes,$\{N^{+,\mathfrak{p}}_{t}\}_{t\in[0,T]}$ and $\{N^{-,\mathfrak{m}}_{t}\}_{t\in[0,T]}$, with controlled stochastic intensities of the form:
@@ 3 Characterization of the value function
$+\left(\lambda^{+}e^{k(s-Z_{+}(y))\Delta^{+}(y)}\sup_{\mathfrak{p}\in\mathbb{R}}e^{-k\mathfrak{p}Z_{+}(y)\Delta^{+}(y)}\left[g(t,y+\Delta^{+}(y),s)-g(t,y,s)+\mathfrak{p}Z_{+}(y)\Delta^{+}(y)\right]\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (3.2)
@@ 6 Conclusion
Although in this paper we focus on a time-horizon that is small enough so that liquidity remains constant, future work will look at the case of non-static liquidity; see e.g., Aqsha et al. (2025); Bergault et al. (2024a). Similarly, one could also look at online learning algorithms (e.g., adaptation and learning with neural contextual bandits or reinforcement learning, as in Duran-Martin et al. (2022); Spooner et al. (2018)) to automate the fees that AMMs charge. Competition between pools is another promising area of future work; see e.g., Boyce et al. (2025); Guo et al. (2024) for studies in limit order books.

### v2

@@ Abstract
Automated Market Makers (AMMs) are emerging as a popular decentralised trading platform. In this work, we determine the optimal dynamic fees in a constant function market maker. We find approximate closed-form solutions to the control problem and study the optimal fee structure. We find that there are two distinct fee regimes: one in which the AMM imposes higher fees to deter arbitrageurs, and another where fees are lowered to increase volatility and attract noise traders. Our results also show that dynamic fees that are linear in inventory and that track changes in the external price are a good approximation of the optimal fee structure and thus constitute suitable candidates when designing fees for AMMs.
@@ 1 Introduction
The early works of Capponi and Jia (2021), Bartoletti et al. (2021), Angeris and Chitra (2020), Angeris et al. (2021), Cartea et al. (2022) and Angeris et al. (2023) first explain how these exchanges operate and introduce the problem of liquidity provision in AMMs. As shown by Cartea et al. (2023, 2024b), Milionis et al. (2022), and Fukasawa et al. (2025), liquidity provision in the absence of fees leads to losses for liquidity providers. There are two well-understood types of losses: loss-versus-rebalancing (LVR) and impermanent loss (IL). Milionis et al. (2024) describe LVR as
the loss incurred by LPs relative to a trading strategy (the “rebalancing strategy”) that holds the same risky positions as the pool, but that trades at market prices rather than AMM prices.
On the other hand, IL is the difference in value between holding the assets inside the pool and holding the assets outside of the pool. The name impermanent is due to the fact that if the market condition reverse, the losses disappear.
To address the IL issue, the literature has generally taken two main approaches. The first is to redesign automated market makers (AMMs) to better account for spot price volatility. For instance, Bergault et al. (2024b) propose an oracle-based AMM, where price discovery is not determined solely by liquidity takers. Instead, their pricing function integrates external information about the current market price of the risky asset.
The second approach focuses on finding trading strategies that allow liquidity providers to adapt to changing market conditions. Capponi and Zhu (2025) tackle the issue of IL by determining an optimal exit strategy for liquidity providers. Fukasawa et al. (2023) address IL in constant function market makers by employing a weighted variance swap for hedging, and in a follow up study, Fukasawa et al. (2024) demonstrate a super-hedging strategy for IL that holds regardless of the oracle price dynamics. Most recently, Sepp et al. (2024) show that IL can be fully hedged using European put and call options.
To find solutions for the loss-versus-rebalancing problem, much of the research has centered on improving AMM designs to make liquidity provision less vulnerable to arbitrageurs. Several notable works in this direction include Adams et al. (2025), who introduce an auction-based AMM in which a pool “manager” captures arbitrage profits and rewards liquidity providers accordingly; Goyal et al. (2023), who convert LPs’ beliefs about future asset values into optimal trading functions; and Cartea et al. (2024a), who generalize constant function market makers into decentralized liquidity pools (DLPs), enabling LPs to mitigate loss-versus-rebalancing by selecting suitable quote and impact functions and by monitoring external information. Another notable contribution is the work by Nadkarni et al. (2024), in which the authors investigate the problem of finding an optimal dynamic trading curve with the focus is specifically on developing a strategy that effectively hedges against loss-versus-rebalancing.
@@
As shown by Milionis et al. (2024), loss-versus-rebalancing can be partially or fully hedged through the collection of fees. Thus, the problem of determining optimal fees for AMMs has been an important issue since the early stages of research in this area. Evans et al. (2021) focused on how to find optimal fees for geometric mean market makers (G3Ms) while Fritsch (2021) studied how to determine optimal fees in a setting of multiple liquidity pools competing with each other for trade volume. He et al. (2024) determined the optimal fee structure for a liquidity provider who allocates their capital between an automated market maker (AMM) and centralized trading venues. However, these works only address the problem of finding constant static fees. Sabate-Vidales and Šiška (2023) address the loss-versus-rebalancing problem by setting dynamic fees although no optimality is established. Hasbrouck et al. (2022) found how higher fees on decentralised exchanges (DEX) can increase trading volume because higher fees attract more liquidity providers and boost the inventory in the DEX. Recently, Aqsha et al. (2025) solve a stochastic leader-follower game between the AMM and the liquidity providers in the venue. They find that under the equilibrium contract, the liquidity providers have incentives to add liquidity to the pool only when higher liquidity on average attracts more noise trading. In their work, the fee that the venue charges is proportional to the trade size.
@@
In this study, we address the problem of determining optimal dynamic fees for AMMs. As argued in Cao et al. (2023), the fees should adjust dynamically in order to reflect market conditions. With the recent introduction of Uniswap v4, the implementation of dynamic fees is now possible. Setting optimal fees is crucial because fees are the primary incentive for liquidity providers to engage with the AMM as they help offset losses from impermanent loss. As a result, a platform that optimizes fee collection can better compensate liquidity providers and, in turn, strengthen its overall market position. Inspired by the work on market making by Avellaneda and Stoikov (2008), [^1] For a comprehensive exposition of the market making problem and its generalisation see the monographs by Cartea et al. (2015) and Gueant (2016). we formulate a stochastic control problem in which the venue aims to maximize the total fees collected in a liquidity pool consisting of two assets: a risky asset $Y$ and a riskless asset $X$. The control processes are $\mathfrak{p}$ and $\mathfrak{m}$, representing the fees charged for selling and buying asset $Y$, respectively. [^2] We choose $\mathfrak{p}$ for plus and $\mathfrak{m}$ for minus because the quantity of asset $Y$ in the pool increases after a trade that pays $\mathfrak{p}$ and decreases after a trade that pays $\mathfrak{m}$. These fees influence the exchange rate faced by liquidity takers: buying $Y$ requires paying more of asset $X$, while selling $Y$ yields less of asset $X$. This, in turn, affects the order flow. We model sell and buy orders using two point processes,$\{N^{+,\mathfrak{p}}_{t}\}_{t\in[0,T]}$ and $\{N^{-,\mathfrak{m}}_{t}\}_{t\in[0,T]}$, with controlled stochastic intensities of the form:
@@ 3 Characterization of the value function
$+\left(\lambda^{+}e^{k(Z_{+}(y)-s)\Delta^{+}(y)}\sup_{\mathfrak{p}\in\mathbb{R}}e^{-k\mathfrak{p}Z_{+}(y)\Delta^{+}(y)}\left[g(t,y+\Delta^{+}(y),s)-g(t,y,s)+\mathfrak{p}Z_{+}(y)\Delta^{+}(y)\right]\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (3.2)
@@ 6 Conclusion
Although in this paper we focus on a time-horizon that is small enough so that liquidity remains constant, future work will look at the case of non-static liquidity; see e.g., Aqsha et al. (2025); Bergault et al. (2024a). Similarly, one could also look at online learning algorithms (e.g., adaptation and learning with neural contextual bandits or reinforcement learning, as in Duran-Martin et al. (2022); Spooner et al. (2018)) to automate the fees that AMMs charge, similar to what has been done in Nadkarni et al. (2023). Competition between pools is another promising area of future work; see e.g., Boyce et al. (2025); Guo et al. (2024) for studies in limit order books.